OpenStudy (anonymous):

Need help please? A box with an open top is to be constructed by cutting a-inch squares from the corners of a rectangular sheet of tin whose length is twice its width. What size sheet will produce a box having a volume of 1248 in3, when a = 3? It wants the width and length.

OpenStudy (anonymous):

I don't know if that will help

OpenStudy (anonymous):

You have to do these problems by drawing them out, look at this pic http://i.ytimg.com/vi/8k7yTz0jYVw/maxresdefault.jpg

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

that really didn't help...

OpenStudy (anonymous):

especially for someone who has no clue how to makeup the formula

OpenStudy (alyssa_xo):

your question didnt copy/paste entirely how big are the cutout squares?

OpenStudy (anonymous):

that's what it looks like.

OpenStudy (alyssa_xo):

oh, okay

OpenStudy (anonymous):

the area is given by 4 little rectangles plus the square in the middle

OpenStudy (anonymous):

the length is x

OpenStudy (anonymous):

find the dimensions of the squares and rectangle

OpenStudy (anonymous):

volume = l * w * h

OpenStudy (anonymous):

okay, if you want me to find the dimensions of the squares and rectangles why are you giving me the formula for volume?

OpenStudy (anonymous):

It's easier for me if we go one baby step at a time and then smush it all together at the end, if you don't mind

OpenStudy (anonymous):

you have a sheet of metal and you want to make a open top box that holds a certain amount of space

OpenStudy (anonymous):

Yes, so lets find the size of each sheet of metal first before me weld them together right?

OpenStudy (anonymous):

for the square on the inside, we would take x-2a multiplied by 2x-2a right?

OpenStudy (anonymous):

sorry im making you do steps you dont need. they mostly what you need in the bottom picture

OpenStudy (anonymous):

x is the length of the sheet of metal, just multiply out all the sides in the bottom pictures and set it equal to the volume they want and then solve for x

OpenStudy (anonymous):

1248 = (a)(2x-2a)(x-2a), a = 36

OpenStudy (anonymous):

Oh, okay. Gotcha, I'm sorry, I'm just trying to understand and I thought the simplest of steps would help me

OpenStudy (anonymous):

ohh so you just take those and put it into to make the volume?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

we should solve for x also correct? so we can find the width and height?

OpenStudy (anonymous):

height = a, length = 2x-2a, width = x-2a though

OpenStudy (anonymous):

a is a constant, 36, so you just need to solve that eqn

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

it says if a is three though? should I plug in 3 or 36?

OpenStudy (anonymous):

oh yea do 3

OpenStudy (anonymous):

okay! one second!

OpenStudy (anonymous):

so far I have 3(2x-6)(x-6) should I distribute three into both parts ?

OpenStudy (anonymous):

or just the first one and then multiply the two?

OpenStudy (anonymous):

I'm sorry for all the stupid questions :/

OpenStudy (anonymous):

you just multiply 1 term out at a time

OpenStudy (alyssa_xo):

just the first one and multiply. if you had something like: \(3[(2x-6)+(x-6) ]\), then you'd distribute into both

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like 4*3*2 what would you do?

OpenStudy (anonymous):

12*2

OpenStudy (anonymous):

same thing with that equation then

OpenStudy (anonymous):

okay, now I have the 6x^2-54x+108

OpenStudy (anonymous):

so I continue to solve for x?

OpenStudy (anonymous):

1248 = (a)(2x-2a)(x-2a), a = 36 416=(2x-3)(x-3) 416=2x^2 -3x - 3x +9 2x^2 - 6x + 9 - 416 = 0 2x^2 - 6x - 407 =0 can you solve from there?

OpenStudy (anonymous):

sorry messed up

OpenStudy (anonymous):

1248 = (a)(2x-2a)(x-2a), a = 36 416=(2x-6)(x-6)

OpenStudy (anonymous):

2x^2 -6x -6x + 36 = 416 2x^2 -6x -6x + 36 - 416 = 0

OpenStudy (anonymous):

2x^2 - 12x -380 = 0

OpenStudy (anonymous):

Now I factor it? I'm really sorry, this is just beyond my knowledge. I think I'm losing my brain because it's late here.

OpenStudy (anonymous):

1248 = (a)(2x-2a)(x-2a), a = 36 should be a = 3 of course

OpenStudy (anonymous):

yeah you can factor or do quadratic formula

OpenStudy (anonymous):

do you understand the steps?

OpenStudy (anonymous):

isn't the quadratic formula easier?

OpenStudy (anonymous):

its what i would do

OpenStudy (anonymous):

Yeah I get the steps, I just don't know why my head isn't freaking functioning lol. I'm really sorry, I'm not this stupid on a daily basis

OpenStudy (anonymous):

dont worry about it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

ugh i keep multiplying wrong, hopefully this is right ill see if i can look it up 1248 = (a)(2x-2a)(x-2a), a = 3 1248 = 3(2x - 6)(x - 6) 2x^2 - 12x - 6x + 36 = 416 2x^2 - 18x - 381 = 0

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-380, not -381

OpenStudy (anonymous):

x = 19, -10

OpenStudy (anonymous):

okay now what?

OpenStudy (anonymous):

but x is the length of the sheet, length cant be negative so dont consider the negative answer

OpenStudy (anonymous):

so 19 should be the answer

OpenStudy (anonymous):

so solve with quadratic formula now?

OpenStudy (anonymous):

19 is one side, then 2*19 is the other

OpenStudy (anonymous):

yeah at 2x^2 - 18x - 381 = 0 you solve with quadratic and should get -10, 19 = x

OpenStudy (anonymous):

-380

OpenStudy (anonymous):

i got to go, do you understand how to finish?

OpenStudy (anonymous):

my laptop is being super slow, sorry

OpenStudy (anonymous):

you cancheck you quadratic formula work with a website

OpenStudy (anonymous):

a is number infront of x^2, b infront of x, c is -380

OpenStudy (anonymous):

then just look at what the problem asks and what the picture showers you just plug into like 2 equations for each side

OpenStudy (anonymous):

the answer was 19 and 38. it wanted the length and width and that was it. thank you sooo soo much

OpenStudy (anonymous):

You have a great day. I'm sorry for the slowness

OpenStudy (anonymous):

no problem